OpenStudy (anonymous):

Given a cylinder with a surface area of 60 units squared and a height of 7 units. Find the volume of a sphere that has the same size radius as the given cylinder. Show all work. Thank You!

OpenStudy (anonymous):

u should know the surface area of a cylinder

OpenStudy (anonymous):

60π units sorry.

OpenStudy (anonymous):

the surface area is ; \[2\pi r^2 + 2\pi rh\]

OpenStudy (anonymous):

how do i find radius

OpenStudy (anonymous):

ok see given that : 2πr^2+2πrh = 60 pi take 2pi common u would get: 2pi(r^2 + h)=60pi r^2+h=30 given that h =7 so r^2+7=30 r^2=23

OpenStudy (anonymous):

then r=√23

OpenStudy (anonymous):

sorry sorry mistake: ok see given that : 2πr^2+2πrh = 60 pi take 2pi common u would get: 2pi(r^2 + rh)=60pi r^2+rh=30 given that h =7 so r^2+7r=30 r(r+7) =30 let r be 3 then 3(3+7) should be equal to 30 NAD IT is so r = 3

OpenStudy (anonymous):

ok now volume is equal to pi r^2 h =63pi unit^3

OpenStudy (anonymous):

have u got my solution tell me if u had got it

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

thanks for giving me medal

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

any more problem

OpenStudy (anonymous):

A solution using Mathematic, with comments is attached. Let me know if you see any errors.

OpenStudy (anonymous):

its 60π units...

OpenStudy (anonymous):

but thank you for your effort.

OpenStudy (anonymous):

it is alright can u tell me frm where u downlaoded mathematica

OpenStudy (anonymous):

www.wolfram.com Mathematic v8, a computer program, is not free, however, they offer the program at a very reasonable price to active students.

OpenStudy (anonymous):

what is the price ?

OpenStudy (anonymous):

I think it is on the order of $125 one time cost as long the student is an actively perusuing a degree or certificate. Check with them on the current situation. I am using the Home version of Mathematic 8 and it cost about $105 to upgrade from version 7 which I believe was $250 at the time of purchase. The Home edition is very restricted and is mainly for hobbyists. I can use it to provide solutions on this site because there is no money or fee involved for my answers.

OpenStudy (anonymous):

but it is available for free download on net

OpenStudy (anonymous):

Yes, they may provide trial copies of the program, however, I believe that the trial copies are programmed to function for around 30 days. After that point in time the program will not function until it is purchased.

OpenStudy (anonymous):

ok